A researcher wishes to estimate the proportion of adults who have​ high-speed Internet access. What size sample should be obtained if she wishes the estimate to be within 0.050.05 with 9090​% confidence if

​a) she uses a previous estimate of 0.36​?

​b) she does not use any prior​ estimates?
Answer #1

Answer

(A) given that

sample proportion p = 0.36

margin of error = 0.05

z critical = 1.645 (using z table)

sample size n = ((1-p)*p)*(z/E)^2

= ((1-0.36)*0.36)*(1.645/0.05)^2

= 0.2304*1082.41

= 249.39

= 250 (rounded to next integer)

(B)

given that

margin of error = 0.05

z critical = 1.645 (using z table)

sample size n = 0.25*(z/E)^2

= 0.25*(1.645/0.05)^2

= 0.25*1082.41

= 270.6

= 271 (rounded to next integer)